OpenStudy (anonymous):

HELP!!! VERY URGENT!!! Consider the differential equation given by dy/dx=xy/3. Find the particular solution y=f(x) to the given differential equation with the initial condition f(0) = 4.

OpenStudy (anonymous):

separate the variables

OpenStudy (whpalmer4):

\[\frac{dy}{dx} = \frac{1}{3}x y\]Separate the variables: \[y^{-1} dy = \frac{1}{3} x dx\]Integrate both sides\[\ln y = \frac{1}{6}x^2 + C\]Take the anti-log of both sides \[y = e^{\frac{1}{6}x^2 + C}\]Now set \(y = 4, x = 0\) and find the value of \(C\).

OpenStudy (anonymous):

before finding C, rewrite it as: \[y = Ce^{\frac{ x^2 }{ 6 }}\]

OpenStudy (amistre64):

need tinier font ... i can almost make it out

OpenStudy (anonymous):

lol

OpenStudy (whpalmer4):

\[\tiny{y=e^{\frac{1}{6}x^2+C}}\] Better? :-)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

much more unreadable, thnx ;)

OpenStudy (whpalmer4):

you're welcome! :-)

OpenStudy (anonymous):

Thank you so much everyone! i really appreciate the help! However, should my final answer be \[y=e^{\frac{ 1 }{ 6 }x^{2}+C}\]

OpenStudy (anonymous):

I really appreciate the help!*

OpenStudy (whpalmer4):

close: you need to find the value of C that makes y = 4 when x = 0

OpenStudy (whpalmer4):

as suggested by @euler271, it may be easier if you rewrite as \[\large{ y = C e^{\frac{1}{6}x^2}}\] before plugging in the initial conditions.

OpenStudy (whpalmer4):

you can do it my way, too, it just takes an extra step to combine the resulting exponent into a nicer form.

OpenStudy (whpalmer4):

I made the font larger this time for @amistre64 but he/she's already left :-)

OpenStudy (anonymous):

I'm kind of stuck trying to solver for \[C\]

OpenStudy (amistre64):

\[\Huge \color{#05ad00}{ y = C e^{\frac{1}{6}x^2}}\]

OpenStudy (whpalmer4):

\[\huge{4 = C e^{\frac{1}{6}(0)^2}}\]\[\huge{4 = Ce^0}\]

OpenStudy (amistre64):

\[\Huge \color{#05ad00}{ y = C e^{\frac{1}{6}x^2}}~:~x=0, y=4\] \[\Huge { 4 = C e^{\frac{1}{6}0^2}}\] \[\Huge { 4 = C e^{0}}\] \[\Huge { 4 = C}\]

OpenStudy (anonymous):

THANK YOU SO MUCH EVERYONE! I REALLY APPRECIATE IT!

OpenStudy (whpalmer4):

Had you not rewritten before finding C, you would get a different value for C: \[\large{4 = e^{\frac{1}{6}(0)^2+C}}\]\[4 = e^C\]\[C = \ln 4\]Then your solution would be \[\large{y = e^{\frac{1}{6}x^2+\ln 4} = e^{\frac{1}{6}x^2}*e^{\ln 4} }\]but of course \[e^{\ln 4} = 4\]so that becomes \[\large{y = 4e^{\frac{1}{6}x^2}}\]just like the other way.